OpenStudy (a1234):

y = -3 cos ((x/3) + (pi/7). Find the period. Is it 3pi?

OpenStudy (anonymous):

For a function of a wave, \[f(x)=A\cos(Bx+C)\] The period of the wave is given \[T=\frac{2\pi}{B}\]

OpenStudy (anonymous):

Think of it as anologous to equation of time-varying displacement of a wave \[y(t)=A\cos(\omega t+\alpha)\] \[\omega\] is the angular frequency, and we know that \[\omega=\frac{2\pi}{T}\]\[\implies T=\frac{2\pi}{\omega}\]

OpenStudy (a1234):

But what is w in this case?

OpenStudy (anonymous):

\[f(x)=-3\cos(\frac{1}{3}x+\frac{\pi}{7})\] Compare it to \[f(x)=A\cos(Bx+C)\] What do you think your B is?

OpenStudy (a1234):

1/3?

OpenStudy (anonymous):

yep! now your job is as simple as finding \[T=\frac{2\pi}{B}\]

OpenStudy (a1234):

That would be 6pi

OpenStudy (anonymous):

Absolutely, solved!

OpenStudy (a1234):

Thanks!